Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 35
Entry time: 24 Apr 2019, 10:05
|
Author: |
Gerd Wellenreuther |
Location: |
General |
Severity: |
Information |
Subject: |
Minutes of the Technical integration meeting 24.04.19 |
|
|
SPB/SFX
- SPB/SFX is preparing for the upcoming beamtime.
- CRL motors being tested. --> Still ongoing.
- CSS interlock tests being planned for tomorrow. --> Still ongoing.
- AE needs to know when Loop 9 and 10 are expected to run. SPB to check.
FXE
- AE needs to get information on PLCs loop splitting. FXE to comment.
- Loop 6 is required to drive 2 servo motors. BB will try to bring this into operation this week. NC checked the setup.
SQS
- AE still needs priorities for the next loops. SQS to comment.
SCS
- Discussion with EET ongoing concerning the installation of cables for end stations. Space problem. --> This is a long term issue, to be discussed in the other meetings, ongoing.
- Commissioning of the motion is ongoing. JT is fixing particular motors / issues.
- Further testing is including interlocks, vacuum and chiller.
- Loop 12 is supposed to be up & running by the end of next week
MID
- For Loop 9 CY is preparing a change request. AE involved. --> still pending
- Loop 12 will be operable this week --> an 100V input is needed, but not readily available. To be checked by AE.
- Priorities for next loops needed by AE to plan the work. --> Still required.
HED
- Loop 7 being commissioned --> still being commissioned, encoders are misbehaving / behaving strange.
- Loop9 (Smaract stages) has to be brough up. --> will be done this week
- Next loops are 8 followed by the priority-items from loop 6.
- Loop10 and 11 regarding PPL will be needed in June. HED to confirm. --> Feedback from MN required.
OPT Lasers
- AE support is needed in order to bring the SCS laser up by the end of may. Mail has been sent. NJardon checks. --> Loop extended and running by the end of next week.
XRO
- Motion loop for Hirex 2 has been delayed until June
XPD
- Issue repoeted from CAS: Fast ADC for MCP in SASE3 duplicates data in the 3 % of cases. DAQ is not responsible. Tests have to be made and for that the ADC servers have to undergo to short interruptions. Will be coordinated with XPD. Should not be e problem. CAS to follow the issue up. --> Unclear, CAS not attending.
VAC
DET
AE
ITDM
- New DACHS-area "Balcony rooms" will be setup. Requires local training by ITDM, which is currently being organized.
EET
- Working on MID SDL and SCS day 2 components.
CAS
SRP
- Schedule changes for the Summer shutdown due to modification of the interlock systems. Additional meetings needed. No info is expected before the beginning of May. --> an uncoordinated email was sent around by MDressel (MPS). SRP will take care of clarifying with MPS what has to be done, when, and will communicate this.
TS
AoB
|